OpenStudy (sharkfinsoup):

Which description means the same as this limit expression? lim f(x)= -infinity x-->infinity A.The graph falls on the left side. B.The graph rises on the right side. C.The graph falls on the right side. D.The graph rises on the left side.

OpenStudy (sharkfinsoup):

I think its A because the f(x)= -infinity mean the graph will go down but the x-->infinity means it will go up on the right

OpenStudy (irishboy123):

i think he means this: \(\Large \lim\limits_{x \to \infty} f(x) = -\infty\) so think about C maybe?

OpenStudy (sharkfinsoup):

why C?

OpenStudy (sharkfinsoup):

@IrishBoy123

OpenStudy (welshfella):

positive values of x are on the right side of the exes

OpenStudy (sharkfinsoup):

ok what does that mean in terms of the question? Are you saying it is going at an upward direction?

OpenStudy (welshfella):

No - it f(x) decends as x in creases to the right

OpenStudy (sharkfinsoup):

Oh so then its B

OpenStudy (sharkfinsoup):

oh that says decend I'm sorry. So it is C

OpenStudy (sharkfinsoup):

@welshfella

OpenStudy (welshfella):

yes

OpenStudy (sharkfinsoup):

Thank you so much!

OpenStudy (irishboy123):

:)